# SWAA: Sliding Window Attention Adaptation for Efficient Long-Context LLMs Without Pretraining

### v4

@@ Abstract
The quadratic complexity of self-attention in Transformer-based Large Language Models (LLMs) renders long-context inference prohibitively expensive. While Sliding Window Attention (SWA), the simplest sparse attention pattern, offers a linear-complexity alternative, naively applying it to models pretrained with Full Attention (FA) causes catastrophic long-context performance collapse due to the training-inference mismatch. To address this, we propose Sliding Window Attention Adaptation (SWAA), a plug-and-play toolkit of recipes that adapt FA models to SWA without costly pretraining. SWAA systematically combines five strategies: (1) applying SWA only during prefilling; (2) preserving “sink” tokens; (3) interleaving FA/SWA layers; (4) chain-of-thought (CoT); and (5) fine-tuning. Our experiments demonstrate that while individual methods are insufficient, specific synergistic combinations can effectively recover original long-context capabilities. After further analyzing performance-efficiency trade-offs, we identify recommended SWAA configurations for diverse scenarios, which achieve 30% to 100% speedups for long-context LLM inference with acceptable quality loss. Our code is available at github.
SWAA: Sliding Window Attention Adaptation for Efficient Long-Context LLMs Without Pretraining
@@ 1 Introduction
Transformer-based Large Language Models (LLMs) Vaswani et al. (2017) demonstrate remarkable capabilities, but their self-attention scales quadratically with the input sequence length, making long context processing inefficient. Sliding Window Attention (SWA), the most straightforward and widely adopted sparse attention pattern, which restricts each token’s attention to a fixed-size local window, reduces the computational complexity to linearity, along with some other benefits (see Appendix A).
Despite its simplicity and ease of use, significant challenges remain in leveraging SWA to accelerate LLM inference. On one hand, due the performance degradation brought by training-inference attention pattern mismatch, to apply SWA to LLMs, most typical solutions involve training a model with SWA from scratch, but they are prohibitively costly and cannot match the performance of state-of-the-art full-causal-attention models like Qwen3 Team (2025b), mainly due to the inability to reproduce pretraining data. On the other hand, training-free methods like streaming attention Xiao et al. (2024) can stabilize LLM outputs by retaining “sink tokens” while applying SWA, which greatly improve efficiency, but inevitably suffer from severe long-context performance degradation possibly due to the inaccessibility of distant tokens’ information Xiao (2025). This motivates a critical, unexplored question: Can we adapt a full-attention model to sliding window attention at low cost while maintaining long-context performance?
We answer Yes to this question by proposing Sliding Window Attention Adaptation(SWAA), a set of recipes for adapting FA-pretrained models to SWA, which requires neither costly pretraining nor modifications to the standard Transformer architecture. Specifically, it systematically combines five practical and composable methods:
@@
    Full Attention (FA) Decode: applying SWA only during the prefilling stage while switching back to full attention for decoding.
@@
    Keep First $k$ Tokens: explicitly preserving attention to the first $k$ “sink” tokens.
@@
    Interleaving FA/SWA layers: mixing full-attention and SWA layers (e.g., applying SWA to half of the layers).
@@
    Chain-of-Thought (CoT): enforcing an explicit "thinking" process during decoding.
  5. 5.
    Fine-tuning with SWA: lightweight SWA-aware supervised fine-tuning on long-context data.
Among these, FA Decode is a novel method we introduce. Keep First $k$ Tokens and FA/SWA Interleaving have been proven effective in prior work Xiao et al. (2024); Team (2024a); Zhang et al. (2024), while CoT and fine-tuning are common LLM techniques. However, the optimal strategy for combining these methods to achieve effective SWA adaptation remains unexplored.
Therefore, in our experiments, we evaluate SWAA on Qwen3 Team (2025b) and Llama3.1 Team (2024b) across several long-context benchmarks, measuring both performance and efficiency under a wide range of SWAA configurations. First, we find that each method makes a distinct contribution, but no single ingredient suffices to make SWA competitive with full attention. Second, we show that specific synergistic combinations of methods can recover a large fraction of the original long-context performance. Third, we analyze the performance-efficiency trade-offs of different SWAA recipes and identify some recommended configurations suitable for different deployment scenarios, which can achieving nearly 100% speedup with 90% accuracy retention in efficiency-first scenarios, or 30% speedup with nearly 100% accuracy retention in quality-first scenarios. This shows SWAA is a flexible toolkit beyond a fixed strategy.
@@
  1. 1.
    We perform the first systematic study on adapting FA-pretrained LLMs to SWA without pretraining, revealing novel insights about how SWA impacts LLMs and providing a foundation for future research in efficient sparse attention.
  2. 2.
    We propose SWAA, a set of practical SWA adaptation recipes that offer a robust performance-efficiency balance for various use cases, accelerating LLM inference from the bottom level.
  3. 3.
    We implement our methods with Flash-Attention Dao (2024) and vLLM Kwon et al. (2023), making it plug-and-play and user-friendly for practical deployment.
@@ 2 Related Works
The $O(N^{2})$ complexity of self-attention in Transformers Vaswani et al. (2017) has spurred a wide field of research about more efficient language model architectures. Among the two most popular technological routes are sparse attention and linear attention.
@@ 2.1 Sparse Attention
Our work falls in this category. Sliding Window Attention (SWA) represents the most basic form of local sparse attention, yet its performance is inherently limited. Therefore, model architectures such as Longformer Beltagy et al. (2020), BigBird Zaheer et al. (2020), and RATTENTION Wang et al. (2025) combine local SWA on most tokens with special global attention on specific tokens to create a more powerful, albeit still sparse, pattern. Popular LLMs like Gemma2 Team (2024a) adopt SWA in half of their layers to balance the efficiency of SWA and performance of FA. Sliding Window Attention Training (SWAT) Fu et al. (2025b) introduces architectural changes, such as sigmoid activation and balanced position embeddings, to stabilize SWA performance. More advanced methods like Deepseek-sparse-attention Yuan et al. (2025); DeepSeek-AI (2025b), although achieving excellent quality, involve more complicated implementation and optimization due to semantic-aware attention operations (e.g., selecting the most important tokens based on attention weights). Nevertheless, nearly all aforementioned methods require pretraining with a specific sparse pattern, which is costly and fails to leverage the advantages of existing pretrained models.
LightTransfer Zhang et al. (2024) is a promising attempt at adapting existing models to SWA without pretraining, which has the same motivation as ours. But it may generalize poorly across model families (see Appendix F).
@@ 2.2 Linear Attention
An alternative approach involves reformulating the attention mechanism entirely to achieve linear, $O(N)$, complexity. This includes methods such as RNN-like linear attention transformers Katharopoulos et al. (2020); Peng et al. (2023); Sun et al. (2023) and structured state-space models (SSMs) like Mamba Gu and Dao (2023). Many works such as Jamba and Nemotron-FlashLieber et al. (2024); Linsong Chu and others (2024); Team et al. (2025); Fu et al. (2025a) interleave linear attention layers with traditional attention layers to create hybrid model structures. While promising, they represent a fundamental architectural departure from the standard Transformer, and their performance is generally weaker than traditional Transformer-based LLMs.
3 Candidate Methods for SWA Adaptation
As previously established, naive application of SWA leads to severe long-context performance degradation. Therefore, we investigate five methods that can potentially facilitate SWA adaptation from distinct perspectives. However, every method except fine-tuning has some drawbacks to LLM inference efficiency, as discussed in Appendix A. Therefore, although these methods are not mutually exclusive, indiscriminately adopting all of them may not achieve the optimal performance-efficiency trade-off. Instead, we must evaluate various combinations to identify the optimal trade-off.
3.1 Full Attention Decode
This method applies SWA only to the prefilling stage. During the decoding (auto-regressive generation) stage, each token still employs full attention, allowing access to all previous tokens in the context. The resulting attention mask is depicted in Figure 1(a).
This novel approach, first proposed in our study, is inspired by human reading comprehension: humans typically scan a passage casually (prefilling) before thinking deeply to formulate an answer (decoding) for a specific problem. We term this strategy "reading casually, thinking carefully." In our design, the SWA-constrained prefilling stage corresponds to casual reading, while the full-attention decoding stage enables careful thinking. This analogy also suggests that Chain-of-Thought (CoT) during decoding may be particularly beneficial, as extended generation (i.e. "thinking") could compensate for the insufficient contextual information gathered during the prefilling stage.
@@ 3.1.1 Keep First k Tokens
Streaming attention Xiao et al. (2024) demonstrates that models pretrained with full attention allocate a disproportionate amount of attention to the initial tokens ("attention sink"), so it permanently retains the attention to these "sink" tokens while using SWA, which successfully maintains the stability of the attention distribution and the model’s output. Our method is basically the same: as shown in Figure 1(b), any subsequent token can attend to its local window and the initial $k$ tokens.
@@
Crucially, we extend this approach further. Streaming attention operates only at the KV cache level; specifically, the KV cache of the sink tokens is retained during decoding, while the prefilling stage is not modified or accelerated at all. In contrast, we directly customize the Flash-Attention-2 Dao (2024) kernel to implement such attention mask, thereby accelerating prefilling via SWA as well, and eliminating the need to modify KV cache.
3.2 Interleaving Layers
This method retains full attention on a subset of layers while applying SWA to the remainder, providing a simple hybrid mechanism to balance the performance of full attention with the efficiency of pure SWA. A common strategy involves designating one in every $n$ layers to use full attention (e.g., layers 0, 2, 4, …retain full attention, while all others use SWA). For example, Gemma-2 Team (2024a) uses SWA only for layers [1, 3, 5, …], and Gemma-3 Team (2025a) uses SWA only for layers [5, 11, 17, …].
However, for an FA-pretrained model, layers may exhibit distinct behaviors, suggesting it may not be optimal to simply assign SWA to layers [1, 3, 5, …]. Some methods like LightTransfer Zhang et al. (2024) have explored more fine-grained layer selection, but we find it is not consistently superior in practice (see Appendix F). Therefore, for simplicity, we still choose the simplest fixed-interval layer selections in our experiments, such as [0, 2, 4, …] and [1, 3, 5, …].
3.3 Chain-of-Thought
Chain-of-Thought (CoT) Wei et al. (2022) is a widely used technique for improving model accuracy via reasoning. With the advent of "thinking" models, such as DeepSeek-R1 DeepSeek-AI (2025a), CoT has evolved from a prompting strategy to an intrinsic LLM capability. However, whether CoT has a specific impact in SWA scenarios remains uninvestigated. We explore this by comparing a thinking model with its non-thinking variant, e.g., Qwen3-4B-Thinking and Qwen3-4B-Instruct, to verify the effect of CoT on SWA adaptation, which can produce more notable differences compared to simply adding CoT prompting to the same model.
3.4 Fine-tuning
Fine-tuning the model with SWA enabled is the most natural way to mitigate training-inference mismatch. Meanwhile, the training data must be long-context, so that SWA actually works.
@@ 3.4 Fine-tuning
However, most available long-context datasets only contain brief ground-truth answers, lacking the reasoning steps required for fine-tuning a "thinking" model. Since our goal is to restore the model’s original capabilities under SWA rather than teach it new ones, we adopt an approach similar to self-distillation Yang et al. (2024). Specifically, we utilize the original full-attention model to generate new answers for the dataset’s questions, and these generated answers are then filtered for correctness using GPT-5-Mini OpenAI (2025), to construct our training dataset. For each question, we sample 4 answers with temperature 1 to enhance diversity, because we find it is slightly better than generating only one answer.
Table 1: The accuracy (%) of Qwen3-4B on LongMemEval, LongBench-V2, and Ruler (without SFT)
No. Window FA layers Keep FA LongMem LB-V2 Ruler
size first decode Think Non-T Think Non-T Think Non-T
0 Full [] 0 False 73.0 62.0 34.6 35.2 85.6 92.8
1 2k [] 0 False 3.2 11.0 9.4 25.8 0.0 0.0
2 8k [] 0 False 13.2 19.8 15.1 22.1 0.0 0.0
3 2k [] 10 False 16.0 15.6 7.7 25.8 0.0 0.0
4 2k [] 0 True 11.8 14.2 26.2 25.2 0.2 0.0
5 2k [1, 3, 5, …] 0 False 13.4 18.4 12.1 23.5 0.0 0.0
6 8k [] 0 True 26.2 25.0 22.8 25.5 0.0 0.0
7 2k [] 10 True 38.2 20.6 25.8 25.2 1.0 0.0
8 2k [] 100 True 50.0 17.8 24.2 26.5 3.2 0.4
9 2k [] 1000 True 50.0 20.2 23.8 25.2 3.2 0.2
10 2k [0, 2, 4, …] 10 False 17.0 14.8 19.8 29.2 0.0 0.0
11 2k [0, 2, 4, …] 0 True 32.2 26.0 23.8 29.9 1.0 2.0
12 2k [1, 3, 5, …] 10 False 25.8 36.4 21.1 29.9 0.0 0.0
13 2k [1, 3, 5, …] 0 True 59.2 34.8 28.5 26.8 22.8 32.8
14 4k [] 10 True 38.0 24.4 27.9 27.5 1.6 2.2
15 8k [] 10 True 49.2 35.2 36.7 30.2 3.2 6.2
16 2k [0, 2, 4, …] 10 True 36.0 17.2 30.5 30.2 10.0 3.0
17 2k [1, 3, 5, …] 10 True 65.0 53.6 35.4 28.6 59.2 75.8
18 2k [1, 3, 5, …] 100 True 68.8 50.6 35.6 31.8 58.0 78.8
19 2k [1, 5, 9, …] 10 True 53.2 31.4 29.6 26.4 9.8 26.4
20 2k [1, 9, 17, …] 10 True 36.4 18.8 28.6 25.4 0.6 0.2
21 2k [3, 7, 11, …] 10 True 54.2 34.6 27.7 26.0 1.6 4.2
22 4k [1, 3, 5, …] 100 True 73.0 54.2 33.4 31.2 65.4 83.6
23 8k [1, 3, 5, …] 100 True 71.6 56.6 35.4 31.5 67.2 86.0
Table 2: The accuracy (%) of Qwen3-4B on LongMemEval, LongBench-V2, and Ruler (with SFT)
No. Window FA layers Keep FA LongMem LB-V2 Ruler
size first decode Think Non-T Think Non-T Think Non-T
0 Full [] 0 False 74.6 63.4 37.9 34.9 88.2 90.6
1 2k [] 0 False 18.8 23.8 7.4 30.9 0.0 0.0
2 2k [] 100 False 15.6 $\backslash$ 6.0 $\backslash$ 0.0 $\backslash$
3 2k [] 0 True 57.9 42.0 29.2 30.2 0.4 0.8
4 2k [1, 3, 5, …] 0 False 63.6 54.6 29.5 31.9 59.0 62.2
5 4k [] 0 True 62.6 $\backslash$ 32.9 $\backslash$ 0.2 $\backslash$
6 2k [] 10 True 56.7 $\backslash$ 28.9 29.2 9.2 14.4
7 2k [] 100 True 62.2 42.6 29.2 30.5 13.4 20.4
8 2k [1, 3, 5, …] 0 True 73.2 58.8 38.3 34.6 74.0 85.8
9 2k [0, 2, 4, …] 0 True 66.0 $\backslash$ 31.5 $\backslash$ 8.0 $\backslash$
10 2k [1, 5, 9, …] 0 True 68.8 47.0 32.0 32.2 7.6 28.0
11 2k [1, 3, 5, …] 100 True 73.2 61.4 37.2 33.9 72.8 89.2
4 Experiment Setup
We organize our experiments around three research questions:
RQ1:
Is SWA adaptation feasible without any additional training?
RQ2:
How much does fine-tuning with SWA improve performance?
RQ3:
Which SWAA configurations achieves the optimal performance-efficiency trade-offs?
4.1 Models
Our primary experiments use Qwen3-4B-Thinking and Qwen3-4B-Instruct Team (2025b). The Thinking variant enforces chain-of-thought (CoT) style reasoning, whereas the Instruct variant usually just answers briefly. To ensure generality, we additionally evaluate Qwen3-30B-A3B-Thinking, Qwen3-30B-A3B-Instruct Team (2025b), and Llama3.1-8B-Instruct Touvron et al. (2023), as shown in Appendix C.

### v5

# SWAA: Sliding Window Attention Adaptation for Efficient and Quality Preserving Long Context Processing

@@ Abstract
The quadratic complexity of self attention in Transformer based LLMs renders long context inference prohibitively expensive. While Sliding Window Attention (SWA), the simplest sparse attention pattern, offers a linear complexity alternative, it suffers from catastrophic long context performance collapse, which stems from two fundamental factors: the training inference mismatch when naively applying SWA to models pretrained with Full Attention (FA), and the inherent structural inability to access distant information when applying SWA to every module at all times. To address these dual challenges, we propose Sliding Window Attention Adaptation (SWAA), a plug and play toolkit of recipes that adapts FA models to SWA without costly pretraining. SWAA systematically combines four core strategies to tackle these distinct issues: (1) Full Attention (FA) Decode and (2) Interleaving FA and SWA layers, which mitigate structural defects by selectively allowing access to distant information; alongside (3) preserving “sink” tokens and (4) lightweight fine tuning, which mitigate the training inference mismatch. Our experiments reveal that while isolated strategies are insufficient, specific synergistic combinations effectively recover long context performance. Despite varying computational overheads, our performance efficiency trade off analysis identifies optimal SWAA configurations for diverse scenarios, achieving 30% to 100% speedups for long context inference with acceptable quality retention. Our code, data and model weights are available at github.
@@ 1 Introduction
Transformer based Large Language Models (LLMs) (Vaswani et al., 2017) demonstrate remarkable capabilities, but their self attention scales quadratically with the input sequence length, making long context processing inefficient. Sliding Window Attention (SWA), the most straightforward and widely adopted sparse attention pattern, restricts each token’s attention to a fixed size local window. This reduces computational complexity to linearity, along with several other benefits (see Appendix A).
Despite its simplicity and ease of use, it causes catastrophic long context performance collapse due to two factors: the training inference mismatch, specifically the model’s parameters cannot adapt to the drift of attention distribution, if directly applying SWA to models pretrained with Full Attention (FA); and the inherent structural difficulty of accessing the information of distant tokens (Xiao, 2025), if SWA is applied to every module at all times, which structurally cuts off all direct information pathways to tokens beyond the local window.
Although various methods that use SWA for efficiency have attempted to address these dual challenges, current solutions struggle to balance flexibility and quality. For instance, training-free methods like streaming attention (Xiao et al., 2024) offer high flexibility, enabling the rapid, plug and play application of SWA on arbitrary FA pretrained models by retaining “sink tokens”. However, they struggle to guarantee long-context generation quality because the structural inaccessibility to distant information still exists. Conversely, models like Gemma2 (Team, 2024a) successfully mitigate this structural limitation by interleaving FA and SWA layers. Yet, this approach necessitates pretraining from scratch with this rigidly defined architecture, which is prohibitively costly, lacks flexibility, and restricts users’ choices of base models.
This dilemma provides our core motivation: Is it possible to leverage the efficiency of SWA to accelerate arbitrary LLMs, while simultaneously resolving these dual challenges at a low cost to maintain acceptable generation quality?
We answer Yes to this question by proposing Sliding Window Attention Adaptation (SWAA), a plug and play toolkit of recipes for adapting FA pretrained models to SWA, which requires neither costly pretraining nor new modules beyond the standard Transformer architecture. Specifically, SWAA systematically combines four practical and composable strategies to tackle these distinct issues.
To mitigate SWA’s structural inability to access distant information, we employ:
@@
    Full Attention (FA) Decode: applying SWA selectively in time (only during the prefilling stage) while switching back to full attention for decoding, allowing the model to retrieve distant context when generating text. This strategy exhibits significant synergy with the use of Chain of Thought (CoT).
@@
    Interleaving FA/SWA layers: applying SWA selectively in space by mixing full attention and SWA layers (e.g., applying SWA to half of the layers), ensuring distant information can still propagate through those FA layers.
To mitigate the training inference mismatch caused by naively introducing SWA, we utilize:
@@
    Keep First $k$ Tokens: explicitly preserving attention to the first $k$ “sink” tokens, to keep the attention distribution stable.
@@
    Fine tuning with SWA: lightweight, SWA aware supervised fine tuning on long context data.
While FA Decode is a novel method we introduce, the other strategies have been proven individually effective in various contexts (Xiao et al., 2024; Team, 2024a; Zhang et al., 2024). However, how combine them for better, cheaper SWA adaptation remains unexplored.
In our experiments, we evaluate SWAA on Qwen3 (Team, 2025b) and Llama3.1 (Team, 2024b) across several long context benchmarks. First, we find that although each strategy is helpful, no single ingredient suffices to make SWA competitive with full attention in answer quality. Second, we show that specific synergistic combinations of these methods can effectively recover long context performance. Third, despite the varying degrees of computational overhead introduced by these methods, our analysis of performance-efficiency trade-offs identifies recommended SWAA configurations for diverse scenarios. For instance, SWAA can achieve nearly 100% speedup with 90% accuracy retention in efficiency first scenarios, or a 30% speedup with nearly 100% accuracy retention in quality first scenarios, making SWAA a flexible toolkit rather than a fixed, rigid strategy.
@@
  • We perform the first systematic study on adapting FA pretrained LLMs to SWA with multiple methods, specifically addressing both the training inference mismatch and the structural defects of SWA, providing a foundation for future research.
  • We propose SWAA, a flexible toolkit of practical SWA adaptation recipes that offer a robust performance efficiency balance for various use cases, accelerating LLM inference from the bottom level.
  • We implement our methods with Flash-Attention (Dao, 2024) and vLLM (Kwon et al., 2023), making SWAA plug and play and user friendly for practical deployment.
@@ 2 Related Works
The $O(N^{2})$ complexity of self attention in Transformers (Vaswani et al., 2017) has spurred a wide field of research about more efficient language model architectures. Among the two most popular technological routes are sparse attention and linear attention.
@@ 2.1 Sparse Attention
Our work falls in this category. Sliding Window Attention (SWA) represents the most basic form of local sparse attention, yet its performance is inherently limited. Therefore, model architectures such as Longformer (Beltagy et al., 2020), BigBird (Zaheer et al., 2020), and RATTENTION (Wang et al., 2025) combine local SWA on most tokens with special global attention on specific tokens to create a more powerful, albeit still sparse, pattern. Popular LLMs like Gemma2 (Team, 2024a) adopt SWA in half of their layers to balance the efficiency of SWA and performance of FA. Sliding Window Attention Training (SWAT) (Fu et al., 2025b) introduces architectural changes, such as sigmoid activation and balanced position embeddings, to stabilize SWA performance. More advanced methods like Native sparse attention (Yuan et al., 2025), although achieving excellent quality, involve more complicated implementation and optimization due to semantic aware attention operations (e.g., selecting the most important tokens based on attention weights).
Nevertheless, nearly all aforementioned methods require pretraining with a specific sparse pattern, which is costly and fails to leverage the advantages of existing pretrained models. Some works explore transforming a full attention trained models to sparse attention through post training or fine tuning. Deepseek sparse attention (DeepSeek-AI, 2025) introduces a lightning indexer and a fine grained token selection mechanism to sparsely process tokens in long context, based on full attention trained Deepseek-V3.1, but still requires nearly 1T tokens for continued pre training. LightTransfer (Zhang et al., 2024) attempts to adapt existing models to SWA without pretraining, which has the same motivation as ours. But it may generalize poorly across model families (see Appendix F).
@@ 2.2 Linear Attention
An alternative approach involves reformulating the attention mechanism entirely to achieve linear, $O(N)$, complexity. This includes methods such as RNN like linear attention transformers (Katharopoulos et al., 2020; Peng et al., 2023; Sun et al., 2023) and structured state space models (SSMs) like Mamba (Gu and Dao, 2023). Many works such as Jamba and Nemotron-Flash (Lieber et al., 2024; Linsong Chu and others, 2024; Team et al., 2025; Fu et al., 2025a) interleave linear attention layers with traditional attention layers to create hybrid model structures. While promising, they represent a fundamental architectural departure from the standard Transformer, and thus must be trained from scratch. Meanwhile, their performance is generally weaker than traditional Transformer based LLMs, and may harm the math and reasoning ability.
3 Method
We incorporate four core adaptation methods into our SWAA framework to address the structural limitations and the training inference mismatch. Their specific operations and motivations are introduced below.
@@ 3 Method
3.1 Mitigating Structural Limitations
3.1.1 Full Attention (FA) Decode (Temporal Selection)
This is a simple yet novel operation, which applies SWA only to the prefilling stage. During the decoding (auto regressive generation) stage, each token reverts to full attention, allowing the model to access all previous tokens across the entire long context. The resulting attention mask is depicted in Figure 1(a).
This approach is inspired by human reading comprehension: humans typically scan a passage casually (prefilling) before thinking deeply to formulate an answer (decoding) for a specific problem. We term this strategy “reading casually, thinking carefully.” In our design, the SWA constrained prefilling stage corresponds to casual reading, while the full attention decoding stage enables careful thinking and unrestricted information retrieval.
Interestingly, this analogy implies that Chain of Thought (CoT) may naturally complements FA Decode, because enforcing an explicit “thinking” process naturally extends the generation length, thus significantly increases the chances and temporal pathways for the model to retrieve or reconstruct distant information that was structurally blocked during the prefilling stage.
3.1.2 Interleaving FA/SWA Layers (Spatial Selection)
This method retains full attention on a subset of layers while applying SWA to the remainder, providing a spatial mechanism to propagate distant information through the network’s depth while balancing the efficiency of pure SWA. A common strategy involves designating one in every $n$ layers to use full attention. For example, Gemma-2 (Team, 2024a) uses SWA only for layers [1, 3, 5, …], and Gemma-3 (Team, 2025a) uses SWA only for layers [5, 11, 17, …].
However, for an FA pretrained model, layers may inherently have distinct behaviors and functions, indicating this method may be sensitive to layer selection strategy. While methods like LightTransfer (Zhang et al., 2024) have explored fine-grained layer selection, we find it is not consistently superior in practice when adapting existing FA models (see Appendix F). Therefore, although the strategy better than fixed-interval selection exists in theory, there is currently no good method to ensure finding it. So, for simplicity and generalizability, we still adopt fixed-interval layer selections in our experiments, such as [0, 2, 4, …] or [1, 3, 5, …].
3.2 Mitigating Training Inference Mismatch
@@ 3.2.1 Keep First $k$ Tokens
Streaming attention (Xiao et al., 2024) demonstrates that models pretrained with full attention allocate a disproportionate amount of attention to the initial tokens (the “attention sink”). Naively applying SWA cuts off access to these sinks, causing severe mismatch. By permanently retaining attention to these initial $k$ tokens while using SWA, the stability of the attention distribution and the model’s output is successfully maintained. As shown in Figure 1(b), any subsequent token can attend to its local window and the initial $k$ tokens.
@@
Crucially, we extend prior approaches. While streaming attention operates only at the KV cache level (retaining the cache of sink tokens during decoding without accelerating prefilling), we directly customize the Flash-Attention-2 (Dao, 2024) kernel to implement this specific attention mask, so as to further accelerate the prefilling stage via SWA and eliminate the need for cumbersome KV cache modifications.
3.2.2 Fine tuning
Fine tuning is a highly direct approach to mitigating the training inference mismatch by updating the model’s weights to adapt to sparse attention patterns. Crucially, fine tuning in our framework does not strictly imply training with naive SWA; rather, it can selectively incorporate any combination of the aforementioned methods. Integrating structural modifications, such as FA Decode, Interleaving Layers, or Keep First $k$ Tokens during the fine tuning phase is relatively straightforward.
@@ 3.2.2 Fine tuning
However, enabling CoT (i.e., fine tuning a “thinking” model) presents a unique data challenge. Most available long context datasets provide only brief ground truth answers, which can be used to fine tune a standard instruct model, but is not applicable to a thinking model since it would severely degrade its reasoning style and capabilities. Since our goal is to restore the model’s original capabilities under SWA configurations rather than teach it new ones, we adopt a self distillation approach (Yang et al., 2024) when fine tuning thinking models. Specifically, we utilize the original full attention thinking model to generate new answers for the dataset’s questions, thereby perfectly preserving the CoT style if it is a thinking model. These generated answers are then filtered for correctness using GPT-5-Mini (OpenAI, 2025) to construct our training dataset. For each question, we sample 4 answers with a temperature of $1.0$ to enhance diversity, which we empirically find yields slightly better adaptation than single answer generation.
4 Experiment
Our experiments are structured to first comprehensively analyze the accuracy of a wide variety of adaptation recipes, covering almost all possibly effective combinations of the components of SWAA, followed by a evaluation of their efficiency, rather than simply evaluating a single configuration where all proposed methods are activated simultaneously. This granular approach is driven by two primary considerations:
First, each structural method introduces varying degrees of computational overhead to LLM inference (as discussed in Appendix A); indiscriminately adopting all of them may get suboptimal efficiency. Instead, we should evaluate multiple configurations to identify the optimal cost-effectiveness.
Second, a systematic ablation by isolating and combining different elements allows us to rigorously verify the necessity and complex dynamics of every component, ultimately providing more comprehensive and useful insights for future research.
4.1 Experiment Setup
4.1.1 Models